Zacharywilly:

The fraction models below represent two fractions of the same whole. How many Fraction 7 over 10s are there in Fraction 1 over 5 ?

mackeulerick2028:

To find how much of 5/8 is you would need to divide 1/2 by 5/8: X = (1/2) / (5/8) When dividing fractions, you reverse the second fraction and multiply: X = (1/2) x (8/5) X = 8/10 Simplify: X = 4/5